# WebRL: Training LLM Web Agents via Self-Evolving Online Curriculum Reinforcement Learning

### v1

@@ Appendix A Training Details
The specific input and output format of WebRL and baselines is shown in Figure 9. The input is composed of task instruction, action history, and HTML of the current page. We process the HTML, simplifying its structure and assigning distinct element IDs to all clickable elements. This facilitates the model’s ability to identify and indicate which specific element requires manipulation. The output specifies the action that needs to be performed on the webpage. The available actions are the same as those defined in WebArena-Lite. The target element for the action is determined by the ”element” argument. To provide more detailed information about the action, we include comments labeled with “# Element:” in the action, which describe the operated element. Similarly, we include comments labeled “# Note:”, which quote relevant information from the current webpage that supports completing the instruction.
ORM: To train the $\mathcal{M}_{\text{ORM}}$, we enhance the WebArena-Lite training dataset by rewriting both the instructions and their associated reward functions. Additionally, we supplement the dataset with examples from WebArena, excluding any cases that are part of the WebArena-Lite test set. Subsequently, we collect rollouts from all baseline methods and WebRL on this augmented dataset, resulting in a total of 12,200 samples. These samples are subsequently used to train the ORM, with the specific hyperparameters listed in Table 5. The prompt for $\mathcal{M}_{\text{ORM}}$ is shown in Figure 18. $\mathcal{M}_{\text{ORM}}$ is required to produce either “YES” or “NO” as its output. To determine the evaluation result, we compare the probabilities assigned to “YES” and “NO” and select the one with the higher probability.
@@
Appendix B Other Quantitative Experiments
@@
Figure 10: Performance variation curves of Llama3.1-8B on each website under WebRL training.
Appendix C Prompts Employed in WebRL
The simple prompt we use to test models including GPT-4-Turbo, GPT-4o, Llama3.1-8B-Instruct, Llama3.1-70B-Instruct, and GLM-4-9B-Chat is shown in Figure 11.
@@ Appendix C Prompts Employed in WebRL
The prompt for generating new instructions is presented in Figure 12.
Figure 12: Prompts for instruction generation.
The prompt used for $\mathcal{M}_{\text{ORM}}$ is shown in Figure 18.
Appendix D Qualitative Examples

### v2

@@ B.1 Details of WebRL
To provide more detailed information about the action, we include comments labeled with “# Element:” in the action, which describe the operated element. Similarly, we include comments labeled “# Note:”, which quote relevant information from the current webpage that supports completing the instruction. An example of the agent’s specific input and output is shown in Figure 11. The input is the observation, while the output specifies the action to be performed on the webpage. The “element” argument identifies the target element for the action. More examples can be found in Appendix E.
The detailed training process of WebRL is shown in Algorithm 1. First, we perform supervised fine-tuning using the WebArena-Lite training dataset. We then initialize the replay buffer and failure set by running the SFT-trained model on instructions corresponding to the WebArena-Lite training dataset. Subsequently, in each phase of the self-evolving curriculum reinforcement learning process, 500 new instructions that meet the filtering criteria are selected from those generated by GPT-4o. Both newly generated interaction data on these instructions and historical data with perplexity between 1/0.95 and 1/0.5 from the replay buffer are used to train the actor and critic. The amount of historical data used is limited to twice the size of the interaction data. The hyperparameters employed in WebRL are presented in Table 5.
B.2 Details of Baselines
For RL-based baselines (except DigiRL), the interaction data from WebRL’s first phase is used, while DigiRL is trained using the first-phase instructions in an online learning setup. Hence, except for DigiRL, the other RL baselines fall under offline reinforcement learning.
We reproduce the same framework used in DigiRL within the WebArena environment. Specifically, we use the same components, including the AWR method, the instruction-level and step-level value functions, and the replay buffer described in DigiRL. To apply DigiRL in WebArena-Lite, the main modifications we make are adjusting the data format to align with WebRL and tweaking certain hyperparameters. DigiRL also conducts 8 rounds of interaction and training. The hyperparameters employed in those baselines are presented in Table 5.
@@ B.2 Details of Baselines
B.3 Details of ORM
WebArena-Lite provides 1,186 training samples each comprising an instruction, a trajectory, and a reward function. To train the $\mathcal{M}_{\text{ORM}}$, we first enhance the WebArena-Lite training dataset by rewriting the instructions. Then, we collect rollouts from all baseline methods on this augmented dataset. These rollouts, combined with the original WebArena-Lite training samples, form the complete training dataset for the $\mathcal{M}_{\text{ORM}}$, with a total of 12,200 samples. We use the associated reward function to label these trajectories, identifying whether they successfully complete the task. These trajectories are subsequently used to train the ORM, with the specific hyperparameters listed in Table 6. The prompt for $\mathcal{M}_{\text{ORM}}$ is shown in Figure 17. $\mathcal{M}_{\text{ORM}}$ is required to produce either “YES” or “NO” as its output. To determine the evaluation result, we compare the probabilities assigned to “YES” and “NO” and select the one with the higher probability. We need to train $\mathcal{M}_{\text{ORM}}$ because new tasks that do not have predefined reward functions will be generated in the subsequent self-evolving curriculum reinforcement learning process. We rely on the trained $\mathcal{M}_{\text{ORM}}$ to judge whether each task is successfully completed.
Algorithm 1 WebRL
Input: WebArena-Lite training set $D_{0}$ and corresponding instruction set $I_{0}$, Base model $\mathcal{M}_{\text{base}}$, Replay buffer $\mathcal{B}$, Failure set $\mathcal{F}$, Reward model $\mathcal{M}_{\text{ORM}}$, Number of phases $N$, Number of instructions per phase $K$
Part 1. SFT Training
1:  $\mathcal{M}_{\text{SFT}}$longleftarrow perform supervised fine-tuning on model $\mathcal{M}_{\text{base}}$ with dataset $D_{0}$
2:  $D_{\text{rollout}}$longleftarrow rollout trajectories of $\mathcal{M}_{\text{SFT}}$ on $I_{0}$
3:  $D_{\text{success}}$, $D_{\text{fail}}$longleftarrow evaluate $D_{\text{rollout}}$ using WebArena-Lite reward function
4:  $\mathcal{B}$longleftarrow $D_{\text{success}}$ Initialize replay buffer with successful trajectories
5:  $\mathcal{F}$longleftarrow $D_{\text{fail}}$ Initialize failure set with instructions of failing trajectoriesPart 2. Self-evolving Curriculum RL
6:  $\pi_{1}$longleftarrow $\mathcal{M}_{\text{SFT}}$ Initialize actor/policy
7:  $V_{1}$longleftarrow $\mathcal{M}_{\text{SFT}}$ with a randomly initialized value head # Initialize critic
8:  for $n$ in 1…$N$ do
9:     $I_{n}$longleftarrow {}
10:     while size($I_{n}$) $\leq$ $K$ do
11:        $I_{\text{generation}}$longleftarrow instructions generated by GPT-4o with instructions from $\mathcal{F}$ as examples # instruction generation process
12:        $I_{\text{filter}}$longleftarrow filter($I_{\text{generation}}$) # instruction filtering process
13:        $I_{n}$longleftarrow $I_{n}\cup I_{\text{filter}}$
14:     end while
15:     $D_{\text{rollout}}$longleftarrow rollout trajectories of $\pi_{n}$ on $I_{n}$
16:     $D_{\text{success}}$, $D_{\text{fail}}$longleftarrow evaluate $D_{\text{rollout}}$ with $\mathcal{M}_{\text{ORM}}$ use $\mathcal{M}_{\text{ORM}}$ to label the rollout trajectories
17:     $D_{\text{experience}}$longleftarrow experiences from $\mathcal{B}$ with perplexity computed by $\pi_{n}$ between $\frac{1}{0.95}$ and $\frac{1}{0.5}$
18:     $\pi_{n+1}$, $V_{n+1}$longleftarrow train($\pi_{n}$, $V_{n}$, $D_{\text{rollout}}\cup D_{\text{experience}}$) # use loss functions from eq. 4 and eq. 7
19:     $\mathcal{B}$longleftarrow $\mathcal{B}\cup D_{\text{success}}$
20:     $\mathcal{F}$longleftarrow $\mathcal{F}\cup D_{\text{fail}}$
21:  end for
@@ B.3 Details of ORM
Appendix C Other Quantitative Experiments
@@
Figure 14 illustrates the error bars for WebRL across each phase. We conduct four sets of experiments, each utilizing a distinct random seed. The results from these four sets of experiments are presented in Figure 14. Despite some fluctuations in accuracy, WebRL consistently demonstrates an overall upward trend. Furthermore, its accuracy consistently surpasses all baseline methods after completing the final phase of training.
Figure 14 presents the success rate of WebRL and DigiRL, both with and without the application of our self-evolving curriculum learning. The results indicate that incorporating curriculum learning significantly enhances the performance of DigiRL. However, its performance still remains below that of WebRL. Conversely, when WebRL is implemented without curriculum learning, its performance experiences a notable decline, yet it still slightly surpasses that of DigiRL.
Figure 16 illustrates the error bars for both the baselines, DigiRL w/ CL, and WebRL. Notably, the performance of WebRL and DigiRL with curriculum learning (CL) shows a higher standard deviation, which can be attributed to the inherent randomness in task generation. Despite this variability, WebRL consistently outperforms all baseline methods, demonstrating its effectiveness.
Figure 16 compares two settings of the replay buffer: one containing only past successful trajectories and the other including all past trajectories, both successful and failed. The results show that including failed trajectories leads to slower and more volatile performance improvement. This issue arises because the advantage of an action in historical trajectories is estimated based on the trajectory’s final reward. Including failed historical trajectories introduces a potential problem: even if certain actions in a failed trajectory are correct, they are likely to receive a negative advantage when the trajectory’s final reward is zero. This reduces the likelihood of selecting these correct actions, thereby hindering effective policy improvement and leading to degraded performance.
Table 7: Performance of different prompts for task generation.
Prompt 1 Prompt 2 Prompt 3
42.4 40.6 43.6
We employ a variety of prompts to assess the stability of the curriculum learning strategy in WebRL. Specifically, we utilize three distinct prompts: Prompt 1 (in Figure 18), Prompt 2 (in Figure 19), and Prompt 3 (in Figure 20). The training results for these prompts, measured after eight phases of training, are presented in Table 7. The results demonstrate that when using a variety of prompts to generate tasks, WebRL consistently delivers strong performance, highlighting the stability of our method.
Figure 12: Performance variation curves of Llama3.1-8B on each website under WebRL training
Figure 13: Error bars of WebRL for each phase.
Figure 14: Comparison of WebRL and DigiRL with and without curriculum learning (CL).
Figure 15: Error bars of baselines, DigiRL w/ CL, and WebRL.
Figure 16: Comparison of WebRL and WebRL with failed trajectories in the replay buffer.
Appendix D Prompts Employed in WebRL
The prompt used for $\mathcal{M}_{\text{ORM}}$ is shown in Figure 17. We require the model to output “YES” or “NO” to determine whether a certain trajectory successfully completes the corresponding task. Considering the limited context size, we only input the HTML content of the last state.
The prompt for generating new instructions is presented in Figure 18. We use tasks that the model fails to complete successfully in previous phases as seeds for generating new tasks of similar difficulty but with greater variety. The prompts shown in Figure 19 and Figure 20 represent alternative versions of the prompts used for task generation.
Figure 18: Prompts for instruction generation.
Figure 19: Prompts for instruction generation (version 2).
Figure 20: Prompts for instruction generation (version 3).
The simple prompt we use to test models including GPT-4-Turbo, GPT-4o, Llama3.1-8B-Instruct, and Llama3.1-70B-Instruct is shown in Figure 21. In this prompt, we define the feasible actions and provide illustrative examples. Additionally, we outline a set of requirements in the “REMEMBER” part to guide the model’s behavior.
@@ Appendix D Prompts Employed in WebRL
The prompt we use to filter the instructions generated by GPT-4o is presented in Figure 22. In this prompt, we provide a rule defining the feasibility of instructions across the five websites in WebArena.
Figure 22: The prompt used to filter instructions.
Appendix E Qualitative Examples